OpenStudy (anonymous):

Rewrite the expression as an algebraic function of x and y. cos(sin−1 x − tan−1 y)

ganeshie8 (ganeshie8):

use angle sum formula

ganeshie8 (ganeshie8):

cos(A-B) = cosAcosB + sinAsinB

OpenStudy (anonymous):

how do I do that?

ganeshie8 (ganeshie8):

\(\cos(\sin^{-1}x - \tan^{-1}y) = \cos(\sin^{-1}x) \cos (\tan^{-1}y) + \sin(\sin^{-1}x) \sin(\tan^{-1}y) \)

ganeshie8 (ganeshie8):

you need to draw two triangles

OpenStudy (anonymous):

why do i need to draw two triangles?

ganeshie8 (ganeshie8):

good question :) cuz u have two different variables

ganeshie8 (ganeshie8):